# Thick isotopy property and the mapping class groups of Heegaard splittings

### v4

@@ 1. Introduction
In [13], Colding, Gabai and Ketover found an effective algorithm to construct the complete list of Heegaard splittings of a non-Haken hyperbolic $3$-manifold. A key of their argument is the thick isotopy lemma ([13, Lemma 2.10]), which allows us to turn the computation of the 00-th homotopy group of the space of Heegaard splittings into a purely combinatorial problem, involving the (crudely) almost normal surface theory. The same strategy is also useful in computing $\pi_{1}(\mathcal{H}(M,\Sigma))$ as stated below. From now on, we fix a Riemannian metric of $M$. Let $\delta>0$. A surface $S$ in $M$ is said to be $\delta$-compressible if there exists a compressing disk $D$ for $S$ such that $\mathrm{diam}\,\partial D\leq\delta$. Otherwise $S$ is said to be $\delta$-locally incompressible.
@@
There have been many efforts to find a finite generating set for the mapping class group of a Heegaard splitting. Possibly the most interesting is $\mathrm{MCG}(S^{3},\Sigma_{g})$, where $(S^{3},\Sigma_{g})$ is a standard genus $g$ Heegaard splitting of the $3$-sphere. It is known that $\mathrm{MCG}(S^{3},\Sigma_{g})$ is finitely generated for $g=2$ by [18] (see also [23]), and for $g=3$ by [17]. However, it is not known if the same is true for $g\geq 4$. On the other hand, a genus $\geq 2$ Heegaard surface in $S^{3}$ is topologically minimal by [2] or [4] (though the disk complex is not finite type). So there might be a good chance to improve our proof to remove the assumption that $M\neq S^{3}$. In fact, much of our argument is still valid when $M=S^{3}$: Lemma 5.5 below is the only place where the assumption $M\neq S^{3}$ is used essentially.
@@
Section 2 is a preliminary towards the proof of Theorem 1.1, including the definition of a crudely almost normal surface. In Section 3, Theorem 1.1 will be proved. Section 4 is a quick introduction to min-max theory. In the final section, we will prove Theorem 1.2 and then conclude with the proof of Corollary 1.3.
Acknowledgement
@@ 2. Normal surface theory
In this section, we recall some definitions and lemmas from [13]. Let $M$ be closed orientable $3$-manifold. Let $\mathcal{T}$ be a triangulation of $M$.
@@
A generic $\mathcal{T}$-isotopy is an isotopy $\{S_{t}\}_{t\in I}$ such that $S_{t}$ is transverse to $\mathcal{T}$ for all $t$ but finitely many points $10<t_{1},\ldots,t_{l}<$ 1. In addition, $S_{{t_{i}}+\epsilon}$ and $S_{{t_{i}}-\epsilon}$ differ by one of the moves shown in Figure 1: When $t$ moves from $t_{i}-\epsilon$ to $t_{i}+\epsilon$, $\Sigma_{t}$ passes either (0) a 00-simplex, (1) a $1$-simplex, (2I) a center tangency with a $2$-simplex or (2II) a saddle tangency. Furthermore, if $|S_{t}\cap\mathcal{T}^{1}|\leq L$ for $t\in I$, $\{S_{t}\}_{t\in I}$ is called a generic $L$-$\mathcal{T}$-isotopy. Finally, for a Riemannian $3$-manifold $M$, an isotopy $\{S_{t}\}_{t\in I}$ is said to be a $C$-isotopy if $\mathrm{Area}(S_{t})\leq C$ for $t\in I$.
@@
[…]
Lemma 2.2 ([13, Lemma 3.2]).
@@
In this section, we prove Theorem 1.1. It is not hard to see the necessity of the theorem. First, recall that a path in $\mathcal{H}(M,\Sigma)$ can be identified with an isotopy of a Heegaard surface. More precisely, as $\mathrm{Diff}(M)\to\mathrm{Diff}(M)/\mathrm{Diff}(M,\Sigma)=\mathcal{H}(M,\Sigma)$ is a fibration [21], any path $\alpha:I\to\mathcal{H}(M,\Sigma)$ lifts to $\tilde{\alpha}:I\to\mathrm{Diff}(M)$ and we can define an isotopy of a Heegaard surface by $\Sigma_{t}:=\tilde{\alpha}(t)(\Sigma)$. Conversely, if an isotopy of a Heegaard surface is given, it defines a path in $\mathcal{H}(M,\Sigma)$ via the isotopy extension theorem. Now if $\pi_{1}(\mathcal{H}(M,\Sigma))$ is finitely generated, we can find a finite collection of isotopies of $\Sigma$ such that any isotopy representing an element of $\pi_{1}(\mathcal{H}(M,\Sigma))$ can be expressed as the product of isotopies in the collection. Thus, $(M,\Sigma)$ satisfies the thick isotopy property.
@@
Let $\{\Sigma_{t}\}_{t\in I}$ be any isotopy with $\Sigma_{0}=\Sigma_{1}=\Sigma$.
@@
$\{\Sigma_{t}\}_{t\in I}$ can be deformed within its homotopy class (as the loop in the space $\mathcal{H}(M,\Sigma)$) so that afterward for all $t$ but finitely many points in $I$, $\Sigma_{t}$ is a crudely almost normal surface with weight bounded above by a universal constant.
This will be proved along the same line as [13, Lemma 3.6]. For our present purpose, an important thing is that the deformation preserves the homotopy class, and this is what we will care about in the proof.
@@
By assumption, we may assume that for $t\in I$ the area of $\Sigma$ is less than $C$ and $\Sigma_{t}$ is $\delta$-locally incompressible. After perturbing $\{\Sigma_{t}\}_{t\in I}$ slightly, $\Sigma_{t}$ is transverse to every simplex of $\mathcal{T}$ for all $t$ but finitely many points $10<t_{1}<\cdots<t_{l}<$ 1. By Lemma 2.2, we may assume that $\Sigma_{t}$ has the weight at most $L:=K(C+1)$ for $t\in I\setminus\{t_{1},\ldots,t_{l}\}$.
For $t\in I\setminus\{t_{1},\ldots,t_{l}\}$, $\Sigma_{t}$ can be turned into a crudely normal surface as follows. Fix a small neighborhood $\mathcal{N}_{t}$ of $\mathcal{T}^{2}$. Note that $\partial{N}_{t}$ consists of $2$-spheres, each corresponding to a $3$-simplex of $\mathcal{T}$. Put $P_{t}:=\partial\mathcal{N}_{t}$. Choose a disk $D\subset P_{t}$ such that $\partial D\subset\Sigma_{t}\cap P_{t}$ and $\mathrm{int}\,D\cap\Sigma_{t}=\emptyset$. Since $\Sigma_{t}$ is $\delta$-locally incompressible and any simplex of $\mathcal{T}$ has diameter at most $\delta$, $\partial D$ must be inessential in $\Sigma_{t}$. So if we compress $\Sigma_{t}$ across $D$, the result is a genus $g$ surface with a $2$-sphere. In other words, the surface is obtained from $\Sigma_{t}$ by a pinch, which is achieved by an isotopy as $M$ is irreducible. If we repeat the above process until all the intersection circles in $\Sigma_{t}\cap P_{t}$ are removed, the resulting surface will be crudely normal with respect to $\mathcal{T}$. Moreover, the weight of the surface is at most $L$ because the normalizing process does not increase the weight.
We need to show that the above normalizing process can be done simultaneously for $t$. The basic idea in the following construction of the isotopy is borrowed from [19].
Let $\{I_{i}\mid 0\leq i\leq l\}$ be a cover of $I$ by closed intervals such that
  • $I_{i}\cap I_{j}\neq\emptyset$ iff $|i-j|\leq 1$,
  • $I_{i}\cap\{t_{1},\ldots,t_{l}\}=\{t_{i}\}$.
Let $\{\tilde{I}_{i}\}$ be a cover of $I$ obtained by enlarging $I_{i}$ slightly. For $i=0,\ldots,l$, choose a neighborhood $\mathcal{N}_{i}$ of $\mathcal{T}^{2}$ so that
  • the boundary, denoted by $P_{i}$, of $\mathcal{N}_{i}$ is transverse to $\Sigma_{t}$ for $t\in\tilde{I}_{i}$,
  • $P_{i}$’s are mutually disjont.
For a simplex $\sigma$ of $\mathcal{T}$, put $\mathcal{N}_{i}^{\sigma}:=\sigma\cap\mathcal{N}_{i}$. By choosing $\mathcal{N}_{i}$ to be small enough, we may assume that any component of $\Sigma_{t}\cap\mathcal{N}_{i}^{\sigma}$ is an annulus, but a single component that is either a disk or a pair of pants, corresponding to the tangency in $\Sigma_{t_{i}}\cap\partial\sigma$.
For $t\in I$, let $\mathcal{C}_{t}$ denote the set of intersecting circles in $\Sigma_{t}\cap\bigcup P_{i}$, where the union is taken for $i$ such that $t\in\tilde{I}_{i}$. For each $c\in\mathcal{C}_{t}$, choose a disk $D$ in $\bigcup P_{i}$ with $\partial D=c$, across which we will compress $\Sigma_{t}$ to remove $c$ from $\Sigma_{t}\cap\bigcup P_{i}$. Denote by $\mathcal{D}_{t}$ the collection of such disks. Pick a family of functions $\{\varphi_{t}:\mathcal{D}_{t}\to[0,2]\mid t\in I\}$ satisfying the following:
  • If $D,D^{\prime}\in\mathcal{D}_{t}$ and $D\subset D^{\prime}$, then $\varphi_{t}(D)<\varphi_{t}(D^{\prime})$.
  • $\varphi_{t}(D)>1$ iff $t\in\tilde{I}_{i}\setminus I_{i}$.
Let $\Gamma$ be the graph in $I\times[0,2]$ given by taking the union of images of $\varphi_{t}$’s. For $t\in I$, the line $t\times[0,2]$ hits $\Gamma$ in finitely many points, say $(t,u_{1}),\ldots,(t,u_{m})$. Going up along $t\times[0,2]$, we find a sequence of disks $\varphi^{-1}_{t}(u_{1}),\ldots,\varphi^{-1}_{t}(u_{m})$. Define the isotopy $\{\Sigma_{tu}\}_{u\in[0,2]}$ to be sequence of pinches corresponding to these disks. Note that if $(t,u_{i})$ is a vertex of $\Gamma$, $\varphi^{-1}_{t}(u_{i})$ is a collection of disks. But the definition of $\varphi_{t}$ guarantees that there is no inclusion relation among such disks. This implies that the corresponding pinches are commutative with each other, and thus $\{\Sigma_{tu}\}$ is well-defined around vertices of $\Gamma$. The restriction of $\{\Sigma_{tu}\}$ on $I\times I$ can be taken to be continuous (though $\{\Sigma_{t2}\}$ cannot be continuous for $t$ because $\mathcal{C}_{t}$ is not constant around $\partial\tilde{I}_{i}$). In conclusion, we obtain a path of isotopies $\{\Sigma_{tu}\}$ with $\Sigma_{t0}=\Sigma_{t}$. Moreover, if $t\in I_{i}$, $\Sigma_{t1}$ is obtained from $\Sigma_{t}$ by successive pinches untill all the intersecting circles in $\Sigma_{t}\cap P_{i}$ is removed.
It remains to check that $\Sigma_{t1}$ is a crudely almost normal surface for $t\in I\setminus\{t_{1},\ldots,t_{l}\}$. Let $\sigma$ be any $3$-simplex of $\mathcal{T}$. Recall that $\mathcal{N}_{i}^{\sigma}=\sigma\cap\mathcal{N}_{i}$. Let $S$ be a component of $\Sigma_{t1}\cap\mathcal{N}_{i}^{\sigma}$. By the choice of $\mathcal{N}_{i}$, any component of $\Sigma_{t}\cap\mathcal{N}_{i}^{\sigma}$ is either a disk, an annulus or a pair of pants. As $\Sigma_{t1}$ is the result of compressing $\Sigma_{t}$ across disks in $\bigcup P_{i}$, $S$ must be either a disk or an annulus. If $S$ is a disk, there is nothing to prove. If $S$ is an annulus, it suffices to see that $S$ is unknotted. Note that such an annulus must be the result of compressing a pair of pants, which corresponds to the saddle tangency in $\Sigma_{t_{i}}\cap\partial\sigma$. Thus, in this case $S$ can be $\partial$-compressed to create a disk, which means that $S$ is unknotted. ∎
[…]
@@ 4. The min-max theorem
  • $\mathrm{\Sigma}_{t}$ converges to $\mathrm{\Sigma}_{t_{0}}$ in Hausdorff topology when $t\to t_{0}$.
@@
  • $\Sigma_{t}$ is a closed genus $g$ surface in $M$ if $t\in\mathrm{int}\,X$. On the other hand, if $t\in\partial X$, $\Sigma_{t}$ is a closed surafce of $\mbox{genus}\leq g$ plus finitely many arcs.
@@
    $\psi(\cdot,t)$ is a self-diffeomorphisms of $M$ for $t\in I\times B^{d}$, and
@@
$W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M):=\inf_{\{\Lambda_{t}\}\in\Pi}\sup_{t\in I\times B^{d}}\mathscr{H}^{2}(\Lambda_{t}).$
A segquence $\{\mathrm{\Sigma}_{t}^{i}\}_{t\in X}$ ($i\in\mathbb{N}$) of sweep-outs in $\Pi_{\{\mathrm{\Sigma}_{t}\}}$ is a minimizing sequence if $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)=\lim_{i\to\infty}\sup_{t\in I\times B^{d}}\mathscr{H}^{2}(\mathrm{\Sigma}_{t}^{i})$. Furthermore, a sequence $\{\mathrm{\Sigma}_{t_{i}}^{i}\}_{i\in\mathbb{N}}$ is a min-max sequence if $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)=\lim_{i\to\infty}\mathscr{H}^{2}(\mathrm{\Sigma}_{t_{i}}^{i})$.
Simon-Smith’s min-max theorem is the following theorem. (The following statement can be found in [13] with minor modification, see [12, 15, 22] for the proof and also [13, Appendix] for the multi-parameter case.)
Theorem 4.1 (cf. [13, Theorem 2.1]).
@@
where $\mathscr{O}$ denotes the subcollection of $\Gamma_{i}$ that is orientable and $\mathscr{N}$ denotes those $\Gamma_{i}$ that are nonorientable, and where $g(\Gamma_{i})$ denotes the genus of $\Gamma_{i}$ if it is orientable, and the number of cross-caps that one attaches to a sphere to obtain a homeomorphic surface if $\Gamma_{i}$ nonorientable.
@@
If $M$ is hyperbolic, $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)\leq 4\pi(g-1)$ by [14, Lemma 9.4]. So, we prove the lemma when $M$ is spherical. Theorem 4.1 shows that $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)=\sum_{i=1}^{k}n_{i}\mathrm{Area}(\Gamma_{i})$ for some embedded minimal surfaces $\Gamma_{i}$ ($1\leq i\leq k$). By Frankel’s theorem [16], the min-max limit is in fact connected and we can express the width as $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)=n\mathrm{Area}(\Gamma)$. If $\Gamma$ is non-orientable, its double cover is stable by [25, Theorem 7.2]. But this is impossible because $S^{3}$ with the standard metric and thus its quotient cannot contain a stable minimal surface. So $\Gamma$ must be orientable and again by [25, Theorem 7.2], the multiplicity $n$ must be one. This together with Choi-Schoen’s area bound [11] for a minimal surface in a spherical $3$-manifold implies
@@
Let $\{\varphi_{k}:S^{d-1}\to\Gamma(\Sigma)\mid k=1,\ldots,n\}$ be a collection maps that represents a finite generating set of $\pi_{d-1}(\Gamma(\Sigma))$. Put $B=B^{d}_{1}$. For each $\varphi_{k}$, we consider the sweep-out $\{\Sigma^{k}_{s}\}$ parametrized by $B$ and given below. (The same construction can be found in [3].)
Take a triangulation $\mathcal{K}$ of $S^{d-1}$ for which $\varphi_{k}$ is simplicial. For a vertex $s$ of $\mathcal{K}$, we will denote by $D_{s}$ a compressing disk for $\Sigma$ that represents $\varphi_{k}(s)\in\Gamma(\Sigma)$. Let $\mathcal{K}^{\prime}$ be the barycentric subdivision of $\mathcal{K}$, and let $\widetilde{\mathcal{K}}$ be the cone over $\mathcal{K}^{\prime}$. So, $\widetilde{\mathcal{K}}$ is a triangulation of $B$. For the center 00 of $B$, define $\Sigma^{k}_{0}=\Sigma$. For a vertex $s$ of $\mathcal{K}^{\prime}$, $\Sigma^{k}_{s}$ is defined as follows. Let $\sigma$ be the smallest simplex of $\mathcal{K}$ that contains $s$. Let $s_{0},\ldots,s_{l}\in\mathcal{K}^{0}$ be the vertices of $\mathcal{K}$ that spans $\sigma$. For $0\leq i\leq l$, take a cylinder $N(D_{s_{i}})\cong D_{s_{i}}\times[-1,1]$ in $M$ such that $D_{s_{i}}\times 0=D_{s_{i}}$ and $N(D_{s_{i}})\cap\Sigma=\partial{D_{s_{i}}}\times[-1,1]$. Furthermore, $N(D_{s_{i}})$’s can be chosen to be mutually disjoint. Define $\Sigma^{k}_{s}$ to be the result of shrinking each annulus $\partial D_{s_{i}}\times[-1,1]$ into an arc across $N(D_{s_{i}})$. By definition, if $s,s^{\prime}\in B$ are vertices lying in the same simplex of $\mathcal{K}$, then $N(D_{s})\cap N(D_{s^{\prime}})=\emptyset$. Thus, the above construction can be extended linearly over any simplex of $\widetilde{\mathcal{K}}$, and defining the sweep-out $\{\Sigma^{k}_{s}\}_{s\in B}$.

### v5

@@ 1. Introduction
In [16], Colding, Gabai and Ketover found an effective algorithm to construct the complete list of Heegaard splittings of a non-Haken hyperbolic $3$-manifold. A key of their argument is the thick isotopy lemma ([16, Lemma 2.10]), which allows us to turn the computation of the 00-th homotopy group of the space of Heegaard splittings into a purely combinatorial problem, involving the (crudely) almost normal surface theory. The same strategy is also useful in computing $\pi_{1}(\mathcal{H}(M,\Sigma))$ as stated below. From now on, we fix a Riemannian metric on $M$. Let $\delta>0$. A surface $S$ in $M$ is said to be $\delta$-compressible if there exists a compressing disk $D$ for $S$ such that $\mathrm{diam}\,\partial D\leq\delta$. Otherwise $S$ is said to be $\delta$-locally incompressible.
@@
Here are a few remarks on the theorem. First, we note that the index of $\Sigma$ is $1$ if and only if $\Sigma$ is strongly irreducible, and so there are many examples of Heegaard splittings satisfying the assumption of the theorem. Unfortunately, we do not know if such examples exist when $d>1$. (This question can also be seen as the special case of Question 5.10 in [3].) However, the examples by Campisi-Rathbun [6] possibly satisfy the assumption. Building on the idea of Bachman-Johnson [4], they constructed examples of hyperbolic $3$-manifolds that contain a Heegaard surface with index $d$ for every $d>0$. Indeed, they proved that there exists a retraction from the disk complex to a sphere $P\subset\Gamma(\Sigma)$ of appropriate dimension. It is likely that we can arrange the construction so that such a sphere is in fact a deformation retract.
We also note that the assumption on $\pi_{d-1}(\Gamma(\Sigma))$ in Theorem 1.2 is used only in the proof of Lemma 5.3, and hence it can be replaced with any condition that implies Lemma 5.3. In particular, it would be interesting to search for examples of Heegaard splittings for which Lemma 5.3 holds without the assumption. In Section 6, we see that this is the case for infinitely many examples of Heegaard splittings of $(\mathrm{surface})\times I$. As a consequence, those Heegaard splittings satisfy the thick isotopy property.
@@
In Section 6, we also establish finite generation of the mapping class groups for infinitely many examples of Heegaard splittings of $(\mathrm{surface})\times I$.
There have been many efforts to find a finite generating set for the mapping class group of a Heegaard splitting. Possibly the most interesting is $\mathrm{MCG}(S^{3},\Sigma_{g})$, where $(S^{3},\Sigma_{g})$ is a standard genus $g$ Heegaard splitting of the $3$-sphere. It is known that $\mathrm{MCG}(S^{3},\Sigma_{g})$ is finitely generated for $g=2$ by [21] (see also [30]), and for $g=3$ by [20]. However, it is not known if the same is true for $g\geq 4$. On the other hand, a genus $\geq 2$ Heegaard surface in $S^{3}$ is topologically minimal by [2] or [7]. (But the disk complex is not of finite type. See Appendix A.) So there might be a good chance to improve our proof to remove the assumption that $M\neq S^{3}$. In fact, much of our argument is still valid when $M=S^{3}$: Lemma 5.5 below is the only place where the assumption $M\neq S^{3}$ is used essentially.
@@
Section 2 is a preliminary towards the proof of Theorem 1.1, including the definition of a crudely almost normal surface. Theorem 1.1 is proved in Section 3. Section 4 is a quick introduction to min-max theory. In Section 5, we prove Theorem 1.2 and Corollary 1.3. In Section 6, we prove that infinitely many examples of Heegaard splittings of $({\rm surface})\times I$ satisfy the thick isotopy property, and as a consequence their mapping class groups are finitely generated. In Appendix A, we see that the disk complex of a Heegaard surface of $S^{3}$ is not homotopy equivalent to a finite simplicial complex.
Acknowledgements
@@ 2. Normal surface theory
In this section, we recall some definitions and lemmas from [16]. Let $M$ be a closed orientable $3$-manifold. Let $\mathcal{T}$ be a triangulation of $M$.
@@
A generic $\mathcal{T}$-isotopy is an isotopy $\{S_{t}\}_{t\in I}$ such that $S_{t}$ is transverse to $\mathcal{T}$ for all $t$ but finitely many points $10<t_{1},\ldots,t_{l}<$ 1. In addition, $S_{{t_{i}}+\epsilon}$ and $S_{{t_{i}}-\epsilon}$ differ by one of the moves shown in Figure 1: When $t$ moves from $t_{i}-\epsilon$ to $t_{i}+\epsilon$, $\Sigma_{t}$ passes either (0) a 00-simplex, (1) a $1$-simplex, (2a) a center tangency with a $2$-simplex or (2b) a saddle tangency. Furthermore, if $|S_{t}\cap\mathcal{T}^{1}|\leq L$ for $t\in I$, $\{S_{t}\}_{t\in I}$ is called a generic $L$-$\mathcal{T}$-isotopy. Finally, for a Riemannian $3$-manifold $M$, an isotopy $\{S_{t}\}_{t\in I}$ is said to be a $C$-isotopy if $\mathrm{Area}(S_{t})\leq C$ for $t\in I$.
@@
Figure 1. When $t$ moves from $t_{i}-\epsilon$ to $t_{i}+\epsilon$, $\Sigma_{t}$ passes either (0) a 00-simplex, (1) a $1$-simplex, (2a) a center tangency with a $2$-simplex or (2b) a saddle tangency.
Lemma 2.2 ([16, Lemma 3.2]).
@@
In this section, we prove Theorem 1.1. It is not hard to see the necessity, i.e. forward implication, in the theorem. First, recall that a path in $\mathcal{H}(M,\Sigma)$ can be identified with an isotopy of a Heegaard surface. More precisely, as $\mathrm{Diff}(M)\to\mathrm{Diff}(M)/\mathrm{Diff}(M,\Sigma)=\mathcal{H}(M,\Sigma)$ is a fibration [26], any path $\alpha:I\to\mathcal{H}(M,\Sigma)$ lifts to $\tilde{\alpha}:I\to\mathrm{Diff}(M)$ and we can define an isotopy of a Heegaard surface by $\Sigma_{t}:=\tilde{\alpha}(t)(\Sigma)$. Conversely, if an isotopy of a Heegaard surface is given, it defines a path in $\mathcal{H}(M,\Sigma)$ via the isotopy extension theorem. Now if $\pi_{1}(\mathcal{H}(M,\Sigma))$ is finitely generated, we can find a finite collection of isotopies of $\Sigma$ such that any isotopy representing an element of $\pi_{1}(\mathcal{H}(M,\Sigma))$ can be expressed as the product of isotopies in the collection. Thus, $(M,\Sigma)$ satisfies the thick isotopy property.
@@
Let $\{\Sigma_{t}\}_{t\in I}$ be any isotopy with $\Sigma_{0}=\Sigma_{1}=\Sigma$. By the argument in [16], we can convert $\Sigma_{t}$ to a crudely almost normal surface with respect to $\mathcal{T}$. Here is a very rough sketch of the argument. By assumption, $\mathrm{Area}(\Sigma_{t})\leq C$ and $\Sigma_{t}$ is $\delta$-locally incompressible for $t\in I$. By Lemma 2.2, $\Sigma_{t}$ is transverse to every simplex of $\mathcal{T}$ for all but finitely many points and $\Sigma_{t}$ has weight at most $L:=K(C+1)$. Using the $\delta$-locally incompressibility condition, for every $3$-simplex $\sigma$ of $\mathcal{T}$, we can pinch off and remove non-disk components of $\Sigma_{t}\cap\sigma$. (We note that a subtle situation may occur when $\Sigma_{t}$ passes through a tangency of type (2b). Around such a tangency, we may be forced to allow an unknotted annulus component. See [16, Proof of Lemma 3.6] for more details.) In summary, we have
@@
$\{\Sigma_{t}\}_{t\in I}$ can be deformed within its homotopy class (as the loop in the space $\mathcal{H}(M,\Sigma)$) so that afterward for all $t$ but finitely many points in $I$, $\Sigma_{t}$ is a crudely almost normal surface with weight bounded above by a universal constant $L>0$.
@@
This follows from [16, Lemma 3.6]. ∎
Consider the graph $\mathcal{G}$ such that each vertex of $\mathcal{G}$ corresponds to a normal isotopy class of crudely almost normal surfaces w.r.t. $\mathcal{T}$ with weight at most $L$, and each edge corresponds to one of the moves (0) – (2b) shown in Figure 1 or a pinch. By Lemma 2.1, $\mathcal{G}$ is a finite graph. In particular, $\pi_{1}(\mathcal{G})$ is finitely generated. By Claim 1, the natural homomorphism $\pi_{1}(\mathcal{G})\to\pi_{1}(\mathcal{H}(M,\Sigma))$ is a surjection. Thus, we conclude that $\pi_{1}(\mathcal{H}(M,\Sigma))$ is finitely generated. ∎
@@ 4. The min-max theorem
  • $\mathrm{\Sigma}_{t}$ converges to $\mathrm{\Sigma}_{t_{0}}$ in the Hausdorff topology when $t\to t_{0}$.
@@
  • $\Sigma_{t}$ is a closed genus $g$ surface in $M$ if $t\in\mathrm{int}\,X$. On the other hand, if $t\in\partial X$, $\Sigma_{t}$ is a closed surface of $\mbox{genus}\leq g$ plus finitely many arcs.
@@
    $\psi(\cdot,t)$ is a diffeomorphism of $M$ for $t\in I\times B^{d}$, and
@@
$W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M):=\inf_{\{\Lambda_{t}\}\in\Pi_{\{\mathrm{\Sigma}_{t}\}}}\sup_{t\in I\times B^{d}}\mathscr{H}^{2}(\Lambda_{t}).$
A sequence $\{\mathrm{\Sigma}_{t}^{i}\}_{t\in X}$ ($i\in\mathbb{N}$) of sweep-outs in $\Pi_{\{\mathrm{\Sigma}_{t}\}}$ is a minimizing sequence if $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)=\lim_{i\to\infty}\sup_{t\in I\times B^{d}}\mathscr{H}^{2}(\mathrm{\Sigma}_{t}^{i})$. Furthermore, a sequence $\{\mathrm{\Sigma}_{t_{i}}^{i}\}_{i\in\mathbb{N}}$ is a min-max sequence if $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)=\lim_{i\to\infty}\mathscr{H}^{2}(\mathrm{\Sigma}_{t_{i}}^{i})$.
Simon-Smith’s min-max theorem is the following theorem. (The following statement can be found in [16] with minor modification, see [15, 18, 27] for the proof and also [16, Appendix] for the multi-parameter case.)
Theorem 4.1 (cf. [16, Theorem 2.1]).
@@
where $\mathscr{O}$ denotes the subcollection of $\Gamma_{i}$ that is orientable and $\mathscr{N}$ denotes those $\Gamma_{i}$ that are nonorientable, and where $g(\Gamma_{i})$ denotes the genus of $\Gamma_{i}$ if it is orientable, and the number of cross-caps that one attaches to a sphere to obtain a homeomorphic surface if $\Gamma_{i}$ is nonorientable.
@@
If $M$ is hyperbolic, $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)\leq 4\pi(g-1)$ by [17, Lemma 9.4]. So, we prove the lemma when $M$ is spherical. Theorem 4.1 shows that $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)=\sum_{i=1}^{k}n_{i}\mathrm{Area}(\Gamma_{i})$ for some embedded minimal surfaces $\Gamma_{i}$ ($1\leq i\leq k$). By Frankel’s theorem [19], the min-max limit is in fact connected and we can express the width as $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)=n\mathrm{Area}(\Gamma)$. If $\Gamma$ is non-orientable, its double cover is stable by [32, Theorem 7.2]. But this is impossible because $S^{3}$ with the standard metric and thus its quotient cannot contain a stable minimal surface. So $\Gamma$ must be orientable and again by [32, Theorem 7.2], the multiplicity $n$ must be one. This together with Choi-Schoen’s area bound [14] for a minimal surface in a spherical $3$-manifold implies
@@
Let $\{\varphi_{k}:S^{d-1}\to\Gamma(\Sigma)\mid k=1,\ldots,n\}$ be a collection of maps that represents a finite generating set of $\pi_{d-1}(\Gamma(\Sigma))$. Put $B=B^{d}_{1}$. For each $\varphi_{k}$, we consider the sweep-out $\{\Sigma^{k}_{s}\}$ parametrized by $B$ and given below. (The same construction can be found in [3].)
Take a triangulation $\mathcal{K}$ of $S^{d-1}$ for which $\varphi_{k}$ is simplicial. For each vertex $s$ of $\mathcal{K}$, choose a representative $D_{s}$ of $\varphi_{k}(s)\in\Gamma(\Sigma)$ so that if $s$ and $s^{\prime}$ are in the same simplex of $\mathcal{K}$, $D_{s}$ and $D_{s^{\prime}}$ are disjoint. Let $\widetilde{\mathcal{K}}$ be the cone over $\mathcal{K}$. So, $\widetilde{\mathcal{K}}$ is a triangulation of $B$. Roughly, $\{\Sigma^{k}_{s}\}_{s\in B}$ is defined as follows. For the center 00 of $B$, set $\Sigma^{k}_{0}=\Sigma$. For each vertex $s$ of $\mathcal{K}$, let $\Sigma^{k}_{s}$ be the result of compressing $\Sigma$ across $D_{s}$. Furthermore, for a barycenter $s$ of an $l$-simplex of $\mathcal{K}$ that is spanned by vertices $s_{0},\ldots,s_{l}\in\mathcal{K}^{0}$, define $\Sigma^{k}_{s}$ to be the result of compressing $\Sigma$ simultaneously across $D_{s_{i}}$. For general $s\in B$, we define $\Sigma^{k}_{s}$ as a linear interpolation of the above construction.
Here is a more formal definition. Let $\tilde{\sigma}$ be a $d$-simplex of $\tilde{\mathcal{K}}$ and let $0=s_{0},s_{1},\ldots,s_{d}$ be the vertices of $\tilde{\sigma}$. Denote by $\sigma$ the $(d-1)$-simplex spanned by the vertices $s_{1},\ldots,s_{d}$. For $1\leq i\leq d$, fix a cylinder $N(D_{s_{i}})\cong B^{2}_{1}\times[-1,1]$ in $M$ such that $B^{2}_{1}\times 0=D_{s_{i}}$ and $N(D_{s_{i}})\cap\Sigma=\partial B^{2}_{1}\times[-1,1]$. Furthermore, $N(D_{s_{i}})$ can be chosen to be mutually disjoint. Let $\rho_{\tau}:B^{2}_{1}\times[-1,1]\to B^{2}_{1}\times[-1,1]$ be the map given by
$\rho_{\tau}(x,u)=((1-\tau b(u))x,u)$
for $(x,u)\in B^{2}_{1}\times[-1,1]$ and $\tau\in I$. Here, $b(u)$ is a bump function. Thus, $\rho_{\tau}$ shrinks the annulus $\partial B^{2}_{1}\times[-1+\epsilon,1-\epsilon]$ to an arc $0\times[-1+\epsilon,1-\epsilon]$. Fix an identification $\iota_{s_{i}}:B^{2}_{1}\times[-1,1]\to N(D_{s_{i}})$ for $1\leq i\leq d$, and define the homotopy $\eta^{s_{i}}_{\tau}:M\to M$ to be
$\eta^{s_{i}}_{\tau}:=\iota_{s_{i}}\circ\rho_{\tau}\circ\iota^{-1}_{s_{i}}$
on $N(D_{s_{i}})$ and to be the identity on the complement of a small neighborhood of $N(D_{s_{i}})$. Regard $\tilde{\sigma}$ as a $d$-cube in $\mathbb{R}^{d}$ so that $0=s_{0}$, the $i$th unit vector represents $s_{i}$, and the other corners correspond to the barycenters of the simplices contained in $\sigma$. Then, every point $s\in\tilde{\sigma}$ can be written as $s=\tau_{1}s_{1}+\cdots+\tau_{d}s_{d}$ for $\tau_{1},\ldots,\tau_{d}\in I$. Now we define $\Sigma^{k}_{s}$ as the image of $\Sigma$ by the composition
$\eta^{s_{1}}_{\tau_{1}}\circ\cdots\circ\eta^{s_{d}}_{\tau_{d}}.$
($\eta^{s_{i}}_{\tau_{i}}$’s are commutative with each other since their supports are mutually disjoint.) Since the above construction is consistent on every intersection between adjacent simplices, we obtain the sweep-out $\{\Sigma^{k}_{s}\}_{s\in B}$.